OpenStudy (anonymous):

for all real numbers X such that 3x-4>5 OR -8x-19<21

OpenStudy (anonymous):

3x-4>5 OR -8x-19<21 3 x > 9 OR -8 x < 40 x > 3 0r x > -5

OpenStudy (anonymous):

\[ (3, \infty) \cup (-5, \infty)= (-5, \infty) \]